Commander Report – 12 April 2018

Commander, PHEnOM Gold Crew, MDRS #193

Mission Sol 4 (Earth Day April 26th, 2018)

Its a beautiful feeling to be on Mars. We have made a home away from home here and are starting to make a special bond with this red planet.

We have adapted to this new routine, with our Executive Officer and Health Science Officer setting of on their morning EVA, while Crew Engineer, Crew Geologist and Astronomer and me working around the hab in the greenhouse, musk observatory and on hab maintenance. The plants in the greenhouse are doing great, we harvested fresh greens today to make salad, soup and to add flavor to our pizzas. While our CE has been doing a great job of maintaining the hab and the vehicles, our CGA has been impressive with his work in the observatory.

The afternoon EVA turned into another adventure when we couldn’t find a road mapped by rovers and previous crews, as it was stormed out and there was no trace left of it. We resorted to the map and GPS coordinates, used our navigational skills and tracked it down. Our efforts were paid off when we reached the canyons at Candor Chasma and were treated to beautiful views. We scouted for emergency shelters, but all that terrain is really flat to have any possibility of emergency shelters, unless we hike down into the canyons. We have left signs to mark part of the road that was washed off, so that future crews are able to find it.

We have had a really busy week so far. Our research projects are coming along great and am happy with the progress we are making.
Planning on winding down the next couple of days as the weather is predicted to be extreme. Will spend time on some team trainings, hab projects and planning for the next week. The team is in great health and spirits and looking forward to the next few days of adventures, research and exploration. I am impressed with the camaraderie and team spirit each crew member has exhibited. Proud to be a part of this crew.

Sol Summary Report – 26 April 2018

Sol 4

Summary Title: Broken Glasses and Martian Pizza

Author: David Attig, Crew Engineer

Mission Status: Super Tremendous Nominal.

Sol Activity Summary: The day started off early with quick individual breakfasts in order to prepare for EVA 8 which left at 0730.

The crew feasted on some homemade pizza, with some fresh salad and soup made with our greenhouse produce. We spent time on personal work and research projects before EVA 9.

While EVA 9 was underway, the HAB was vacuumed and progress made on experiments.

Look Ahead Plan: Tomorrow we will stick close to the Hab due to predicted hot weather and high winds in the afternoon. We will have two EVAs to practice on ATVs, and will spend the afternoon in the Hab practicing medical emergency situations and reviewing fire safety.

Anomalies in work:

1. There was a pair of sunglasses lost during the afternoon EVA. They were found, but had been driven over by a Rover.

2. There was a missed picture opportunity; the crew member who was asked to take it did a poor job and the crew member who asked for it did not review the picture before returning from EVA.

Weather: Clear skies in AM, chilly in morning and quickly warming during the day. Warm and breezy in PM.

Crew Physical Status: Nominal

EVA:

EVA 8 – went well, gray moon area was explored. There is incredible scenery, the landscape matches up with its name very well. Upon returning, they brought materials in from the RAM to work on in the HAB.

EVA 9 – Saw a fantastic view of Candor Chasma.

Operations Report – 26 April 2018

SOL: 4

Name of person filing report: David Attig

Non-nominal systems: Com radio #6, Sim suits #10, 2, 5, 4, 6.

Notes on non-nominal systems: Status unchanged since yesterday.

Generator (hours run): 9:00 pm to 9:00 am.

Solar— SOC 80 % (Before generator is run at night)

Diesel – 70 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 5 gallons.

Water (trailer) – 550 gallons.

Water (static) – 305 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134126.7

Toilet tank emptied: yes
Date last emptied: April 24

ATV’s (Honda, 300, 350.1, 350.2, 350.3): No

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 0 Gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: 350.2 is out of service, due to a previous issue.

Deimos rover used: Yes
Hours: 124.8
Beginning charge: 100%
Ending charge: 94 %
Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: Yes
Hours: 35.0
Beginning charge: 100 %
Ending charge: 91 %
Currently charging: Yes

Opportunity rover used: Out Of Service
Hours: 31
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: 91 %
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: I took pans out of broiler underneath
oven; It is an older style gas stove, with a broiler (not a drawer)
underneath the oven. Too many pans there could block airflow for oven
burner.

Summary of internet: As of 9:30 AM, we have 40% of our internet
allowance remaining. The funny thing is that our internet while we
actually do have some percentage remaining doesn’t seem any faster
than at night when we are hard throttled.

Summary of suits and radios: All one piece suits are working (and
were used today). Radios are nominal.

Summary of Hab operations: Supported two (4 hour) EVAs.

Summary of GreenHab operations: Picked some plants for lunch today.

Summary of ScienceDome operations: Samples placed in incubator.

Summary of RAM operations: Morning EVA team brought materials in from
RAM upon returning. New radios are in the RAM airlock for us.

Summary of health and safety issues: None to report.

EVA Report – 26 April 2018

Sol 4 Crew 193 EVA 8&9 Summary

Earth-Date 26APR2018

EVA 8

EVA Gold Crew Members – Gold001 (XO Campbell) / Gold004 (HSO Pandya) (Informally: “Team Canada”)

EVA Objectives – Summary

  1. Exploration of The Moon – Completed

  • Distance covered by foot – 8.13km

  • Steps – 11,000

  1. Scouting and Mapping of Emergency Shelter Sites – Noted / Raw data here

  2. EVA Comfort, Well-Being, Safety and Optimization Study / Raw data here

  3. Visor Fogging Study – Data collected / Raw data here

  4. Moving signs/painting equipment from RAM to airlock for later use outside observatory – done

Issues/Mitigation

  • Turned back after 1h15min due to extreme heat, returned to base 40-50min early to avoid medical issues related to heat exhaustion/hyperthermia, have suggested to EVA 9 crew to cut their EVA short and/or bring water into field

ATVs/Rovers Used

  • Deimos – 100% start; 94% end

  • ATV use on hold until refueled

Safety Measures & Preparation

  • Sunscreen/Sun protection – Yes/Medical Exemptions for UV eye-protection provided

  • Adequate hydration – No

  • Adequate protection – Flightsuit/footwear/visor/glove precautions – Yes

Sol 4 Crew 193 EVA 9 Summary

Earth-Date 26APR2018

EVA 9

EVA Gold Crew Members – Gold000 (CO Patil-Sabale) / Gold002 (CE Attig) / Gold005 (CGA Shear)

EVA Objectives – Summary

  1. Exploration of Candor Chasma – Accomplished

  • Total distance covered – 19.69 km (roundtrip and side trips)

  1. EVA Comfort, Well-Being, Safety and Optimization Study / Raw data here

  2. Visor Fogging Study – Data collected / Raw data here

Issues/Mitigation

  • Missed the split in Cactus Road and went south instead of northeast, have set up a sign at (0519955, 4251570) for future follow ups.

ATVs/Rovers Used

  • Spirit and Curiosity.

Safety Measures & Preparation

  • Sunscreen/Sun protection – Yes

  • Adequate hydration – Yes

  • Adequate protection – Flightsuit/footwear/visor/glove precautions – Yes

Science Report – 26 April 2018

1. Emergency EVAC EVA: No locations were scouted during either EVA today. Planning for an EVA dedicated to emergency procedures is in the works.

2. Sunspot and Solar Flare Monitoring: Our CGA continued his task of monitoring the sun for sunspot activity. The sunspots are getting smaller – they were not visible with the eye but could still be seen with the Skyris camera.

3. Shortwave Texting on EVA’s: Beartooth devices still not working. Hand signal communication was used instead.

4. GPS Route Measurements: GPS was taken on EVA 8 this morning which proved very helpful in re-locating the rover after exiting the moon valley area. A picture was taken of the GPS device when EVA 8 parked for exploration and the crew the followed the GPS back to that set of coordinates.

5. Waterless dish cleaning: No samples were taken today. No growth seen on samples from previous day. Previous day samples were put into the incubator in the science hab today to speed bacterial growth.

6. The MAG (Maximum Absorbency Garment) utilization study: No progress today. Possible that the continuation of this study from Crew 192 will be scrapped.

7. Spacesuit Helmet Fogging: Neither crew experienced fogging of the masks today. Both crews used a defogging spray and non-scratch cloth on the masks prior to their EVAs today. One crew member thought their mask had fogged, but it turned out to just be dirt.

8. Crew Comfort, Health and Safety on EVA: Measurements of the crew’s weight, blood pressure, SpO2 and pulse were taken before and after the EVA. Maximum heart rate was tracked by a mobile monitor during the EVA when the astronaut felt that they were working very hard (normally climbing a large hill). Gold 01 is able to track his heart rate over the entire course of the EVA thanks to mobile technology. Results can be exported and used in this study. Analysis will be done as the mission continues.

Astronomy Report – 27 April 2018

Name: Eric Shear

Crew: 193
Date: April 26, 2018

Sky Conditions: Clear

Wind Conditions: None

Observation Start Time: 0800

Observation End Time: 845

Summary:

Our (Crew Geologist Astronomer) CGA continued his task of monitoring the sun for sunspot activity. The sunspots are getting smaller – they were not visible with the eye but could still be seen with the Skyris camera. The triple-sunspot observed yesterday has now been reduced to two, with the third almost indistinguishable from the surrounding granules.

No prominences, flares, or filaments were seen.

Photos were taken with the Skyris camera and the best image processed in Autostakkert, Registax, Photoshop, and PowerPoint to colorize and bring out contrast. The result is attached.

Objects Viewed: Sun

Problems Encountered: None

GreenHab Report – 27 April 2018

Name: Shawna Pandya

Date: 26 Apr 2018

Environmental control: Ambient with door open but only for few hours in the afternoon

Shade cloth: on

Fan: not used today

Average temperatures: 42.1 C

Low: 26.6℃

High: 42.2℃

Hours of supplemental light: None

Changes to crops: None

Daily water usage for crops: 5 Gallons

Time(s) of watering for crops: Twice: Once at 0900

Morning research observations: None

Changes to research plants: No.

Aquaponics: None

Narrative: Picked kale, thyme, rosemary, sage, swiss chard and red lettuce.

Support/supplies needed: None